# Supertranslation ambiguity in post-Minkowskian expansion

### v2

@@ 2 Self-part at the second order
where $\gamma_{AB}$ is the boundary metric of the celestial sphere and $D_{A}$ is the derivative with respect to $\gamma_{AB}$. The capital index will be lowered and raised by $\gamma_{AB}$ and its inverse metric. Here, $\bar{\beta}(x^{A})$ is an arbitrary function on the celestial sphere which characterizes the first order supertranslation. The effective stress tensor at the second order that are relevant for deriving the second order solution space is given by
$T^{E}_{rr}=-\frac{C^{AB}C_{AB}}{2r^{4}}+{\cal O}(r^{-5}),\quad\quad T^{E}_{rA}=\frac{D_{A}(C^{AB}C_{AB})}{4r^{3}}+{\cal O}(r^{-4}),$ (8)
$T^{E}_{ur}={\cal O}(r^{-4}),\quad\quad T^{E}_{AB}=\frac{\gamma_{AB}C^{EF}N_{EF}}{2r}+{\cal O}(r^{-2}),\quad\quad N_{AB}=\partial_{u}C_{AB},$ (9)
$T^{E}_{uA}=-\frac{D^{B}C_{AE}{N_{B}}^{E}+3C_{AB}D_{E}N^{BE}}{2r^{2}}+{\cal O}(r^{-3}),$ (10)
$T^{E}_{uu}=-\frac{N^{AB}N_{AB}+2C_{AB}\partial_{u}N^{AB}}{2r^{2}}+{\cal O}(r^{-3}).$ (11)
3 Solution space at the second order
@@
We will solve the second order theory in series expansion near future null infinity. All the metric components are assumed to be given in $\frac{1}{r}$ series expansions. We will adapt the Bondi gauge to the linearized theory where it is easy to read the following conditions
$h_{rr}=h_{rA}=0.$ (12)
The determinant condition in the Bondi gauge will induce constraints between the trace of $h_{AB}$ and the data at ${\cal O}(G)$. The precise condition is
@@
where $c_{AB}$ and $d_{AB}$ are traceless. This condition seems artificial. But the determinant condition indeed mixes data between different orders of the expansions in $G$. One can alternatively use the Newman-Unti (NU) gauge [43] (see also the application in linearized gravity [44]) where one can set the gauge condition $h_{ur}=0$. We have checked that transforming the solution in the NU gauge to the Bondi gauge will fix the subleading order of $h_{AB}$ as (13). The boundary conditions adapted from the Bondi gauge are
$h_{ur}={\cal O}(r^{-2}),\quad\quad h_{uA}={\cal O}(1).$ (14)
The effective stress tensor is conserved outside the trajectory of the pointlike bodies, see the proof in Appendix B. We can use the well know organization of Einstein equation in Bondi gauge [3, 4, 45, 44] to solve the linearized theory at the second order near null infinity. The four hypersurface equations $E_{r\alpha}=T^{E}_{r\alpha}$ will determine $h_{ur}$, $h_{uA}$, and $h_{uu}$. In series expansion, we have
@@
$h_{uA}=D^{B}c_{AB}+\frac{1}{r}\left[\frac{2}{3}(n_{A}+uD_{A}m_{b})-\frac{1}{16}D_{A}(C_{AB}C^{AB})\right]+{\cal O}(r^{-2}),$ (16)
@@
where $m_{b}$ and $n_{A}$ are integration constants from the differential equations of $r$. Two standard equations, the traceless part of $E_{AB}=T^{E}_{AB}$, will fix the evolution of data at each order of $h_{AB}$ except $c_{AB}$. In particular $\partial_{u}d_{AB}=0$ from the leading order. The trace part of $E_{AB}=T^{E}_{AB}$ is trivial when the above six equations of motion are satisfied. Three supplementary equations will fix the time evolution of the integration constants as
@@
$\partial_{u}m_{b}=\frac{1}{4}D_{A}D_{B}n^{AB}-\frac{1}{8}N_{AB}N^{AB},\quad\quad n_{AB}=\partial_{u}c_{AB}.$ (19)
The results are in consistent with the solution space of full Einstein theory near null infinity which is detailed in Appendix C. Inserting the precise formula of $C_{AB}$ from previous section, the evolution relations are reduced to
@@
In the next section, we will use the 4-momentum and angular momentum at the next-to-leading order in $G$ to fix $n_{A}$ and $m_{b}$.
4 Four-momentum and angular momentum
The Bondi 4-momentum and angular momentum at the leading order coincide with the ADM definitions [40],
@@ 4 Four-momentum and angular momentum
The next-to-leading order contribution of the Bondi 4-momentum and angular momentum at null infinity are [46, 47]
@@
where $\text{d}\Omega$ is the integral measure on the celestial sphere and $Y^{(i)A}\frac{\partial}{\partial x^{A}}$ are three Killing vectors of the celestial sphere, which are related to the normal vector $n^{i}$ by
$Y^{(i)A}=-\epsilon^{AB}D_{B}n^{i},\quad\quad x^{i}=rn^{i}$ (24)
where $x^{i}$ are the spatial parts of the Cartesian coordinates. Since the self-part second order solution does not include the stress tensor from the pointlike bodies, the physical charges should not be changed, namely the next-to-leading order Bondi 4-momentum and angular momentum should vanish. Those conditions will completely fix $m_{b}$ and $n_{A}$ as
@@
$n_{A}=\frac{1}{4}\int\,\text{d}u\,\left(D^{B}D_{A}D^{E}c_{EB}-D^{2}D^{B}c_{BA}-uD_{A}D_{B}D_{E}n^{BE}\right)+......,$ (26)
where the ellipsis denotes $u$-independent total derivative terms which can not be fixed by the charge analysis. Note that we have applied the identity [48, 29]
$-\frac{1}{4}Y^{(i)A}C_{AB}D_{E}C^{BE}=\frac{1}{4}D^{A}\bigg{[}Y^{(i)B}C_{AB}(D^{2}+2)\left[f(x^{A})-\bar{\beta}(x^{A})\right]\\ +\frac{1}{2}Y^{(i)}_{A}\bigg{(}(D^{2}+2)\left[f(x^{A})-\bar{\beta}(x^{A})\right]\bigg{)}^{2}\bigg{]},$ (27)
according to the expression of $C_{AB}$ in (6). Note also that
$Y^{(i)A}\left(D^{B}D_{A}D^{E}c_{EB}-D^{2}D^{B}c_{BA}\right)=\\ D^{B}\bigg{(}Y^{(i)A}D_{A}D^{E}c_{BE}-D^{A}Y^{(i)}_{B}D^{E}c_{AE}\\ -Y^{(i)A}D_{B}D^{E}c_{AE}+D_{B}Y^{(i)A}D^{E}c_{AE}+2Y^{(i)A}c_{AB}\bigg{)},$ (28)
is a total derivative term when $Y^{(i)A}$ is a Killing vector on the sphere which satisfies
$D_{B}Y_{A}^{(i)}+D_{A}Y_{B}^{(i)}=0,\quad\quad D_{B}D_{C}Y_{A}^{(i)}=\gamma_{AB}Y_{C}^{(i)}-\gamma_{BC}Y_{A}^{(i)}.$ (29)
Though we have fixed $m_{b}$ and $n_{A}$ from the conserved quantities, the relevance to the first order supertranslation is not yet uncovered. We will compute the transformation law of $m_{b}$ and $n_{A}$ under the first order supertranslation and prove that a generic first order supertranslation can not preserve the form of $n_{A}$ in (26).
5 First order supertranslations
A supertranslation connecting two solutions in Bondi gauge at the linear order is characterized by the transformation from $(u^{\prime},r^{\prime},{x^{\prime}}^{A})$ to $(u,r,x^{A})$,
$u^{\prime}=u-GT(x^{A}),\quad\quad r^{\prime}=r-\frac{G}{2}D^{2}T,\quad\quad{x^{\prime}}^{A}=x^{A}+\frac{G}{r}D^{A}T.$ (30)
However, this will not preserve the new metric
$g_{\mu\nu}={g^{\prime}}_{\alpha\beta}\frac{{\partial}{x^{\prime}}^{\alpha}}{\partial x^{\mu}}\frac{{\partial}{x^{\prime}}^{\beta}}{\partial{x}^{\nu}},$ (31)
in Bondi gauge at ${\cal O}(G^{2})$. Considering the corrections from ${\cal O}(G^{2})$, the coordinates transformation that preserves the Bondi gauge up to relevant order in $\frac{1}{r}$ expansion is
@@ 5 First order supertranslations
$\quad\quad\quad+2D_{A}TD^{A}T-\frac{1}{2}D^{2}TD^{2}T\bigg{)}+\frac{G^{2}}{6r^{2}}D^{AB}D_{A}D_{B}T$ (33)
@@
$\quad\quad\quad-\gamma^{AB}\partial_{B}D^{C}TD_{C}T+\partial_{C}\gamma^{AB}D_{B}TD^{C}T\bigg{)}-\frac{G^{2}}{3r^{3}}D^{AB}D_{B}T+{\cal O}(G^{3}),$ (34)
where $\gamma=\sqrt{\det{\gamma_{AB}}}$ and $D_{AB}$ is the subleading order term of $H_{AB}$ in asymptotic expansion. We have checked that $D_{AB}$ will not affect the transformation law. So we will not give its precise form. Note that we will check the transformation laws under the first order supertranslation. The second order supertranslation is not considered. For computational simplicity, we choose the complex stereographic coordinates $A=(z,\bar{z})$ on the celestial sphere. The transformation laws of the Bondi mass aspect, angular momentum aspect, and asymptotic shear at the linear order is obtained as
$M_{b}={M^{\prime}}_{b},\quad\quad N_{A}={N^{\prime}}_{A},\quad\quad C_{AB}={C^{\prime}}_{AB}+(2D_{A}D_{B}-\gamma_{AB}D^{2})T,$ (35)
where the primed quantities are functions in the old coordinates system. The results are in accordance with the fact that the leading order Bondi 4-momentum and angular momentum are irrelevant to the first order supertranslation [40]. The transformation laws at the second order are
@@
$n_{A}={n^{\prime}}_{A}-D_{A}(T{M^{\prime}}_{b})-2{M^{\prime}}_{b}D_{A}T.$ (37)
The results are consistent with the finite BMS transformation results in [49, 50, 51] when the linear order metric is substituted. Note that we have used the expression of $C_{AB}$ in (6) to verify that
$D^{\bar{z}^{\prime}}D^{\bar{z}^{\prime}}{C^{\prime}}_{\bar{z}^{\prime}\bar{z}^{\prime}}-D^{z^{\prime}}D^{z^{\prime}}{C^{\prime}}_{z^{\prime}z^{\prime}}=0,$ (38)
@@
when deriving (37). Clearly, the condition of preserving $n_{A}$ in the form of (26) in the new coordinates system will reduce the supertranslation to one translation along time direction $u$, considering that $M_{b}=-\sum{\frac{m}{(v_{\mu}n^{\mu})^{3}}}$ can not be a constant for arbitrary numbers of pointlike bodies. Even for the case of a single particle, $M_{b}$ is not necessary to be a constant and a supertranslation can change the angular momentum of the system which agrees with that the boost charges of the supertranslated Schwarzschild black hole are generically non-vanishing [52]. Note that the velocity of the particle essentially is a boost of the Schwarzschild solution [53].
6 Concluding remarks
@@
In this paper, we apply the charge conservation to prove that the first order supertranslation should be ruled out in the PM expansion which can serve as the foundation of uniquely determining the angular momentum loss in gravitational scattering. It is worthwhile to point out that ruling out the first order supertranslation only means that the first order supertranslation parameter is fixed, but not necessary to be zero. The precise formula of the first order supertranslation parameter can be obtained for a precise gravitational scattering, such as gravitational Bremsstrahlung [54, 41, 10]. It is amusing to notice that once the first order supertranslation is fixed, the complete first order metric is fixed. The derivation of the other parts of the second order metric are relevant to the first order metric, but can not affect the first order metric. Then the key point of our prescription is that we use part of the higher order solutions in $G$ expansion to fix the lower order supertranslation. It is interesting to see if such idea can be valid at higher order in $G$ to fix the complete supertranslation freedom.
To close this paper, we will comment on the physical consequence of the first order supertranslation in the PM expansion. In the infinitesimal version, a diffeomorphism transformation can be expanded in $G$ as
${\cal L}_{\xi}g_{\mu\nu}={\cal L}_{\xi}(\eta_{\mu\nu}+GH_{\mu\nu}+G^{2}h_{\mu\nu}).$ (39)
If one expands the symmetry parameter as $\xi=G\xi^{(1)}+G^{2}\xi^{(2)}$, the induced first order infinitesimal gauge transformation is
$\delta H_{\mu\nu}={\cal L}_{\xi^{(1)}}\eta_{\mu\nu}.$ (40)
Note that this transformation only changes the perturbative metric $H_{\mu\nu}$. We do not consider its action on the background spacetime. So it does not change the frame of the observer and it is a pure gauge transformation of the perturbative metric. At the second order, the induced gauge transformation is
The first piece on the right hand side is the independent second order gauge transformation. The second piece is induced from the first order gauge transformation. It is not a pure gauge transformation at the second order because it will change the first order metric. One can effectively consider that the second order perturbative theory is living on a curved background with metric $\eta_{\mu\nu}+GH_{\mu\nu}$. The effective stress tensor coupled to the second order theory is from curved effect of the background spacetime. Then the first order gauge transformation does change the spacetime metric, hence change the frame of the observer. Note that the translations along spatial dimensions correspond to the shift of the origin of the spacetime which has also been ruled out as they will change the angular momentum. Thus, a first order supertranslation (except the translation along $u$) will set the second order theory to a different frame or a different origin choice than the first order theory. Such interpretation is in consistent with the full Einstein theory [36] where memory effect will change the frame of observer which leads to the supertranslation ambiguity of angular momentum. In the PM expansion, there is no dissipation from the self-part at the second order. The charge conservation can serve as calibration of frames and origin choices of the two observers for describing the first order and second order perturbative theories. The fact that theories at different orders can be in different frames or with different origin choices reflects the non-linear effect from interactions between the gravitational fields and the observers in the full Einstein theory.

### v3

@@ 2 The resolution to supertranslation ambiguity
where $\gamma_{AB}$ is the boundary metric of the celestial sphere and $D_{A}$ is the covariant derivative with respect to $\gamma_{AB}$. The capital index will be lowered and raised by $\gamma_{AB}$ and its inverse metric. Here, $\bar{\beta}(x^{A})$ is an arbitrary function on the celestial sphere that characterizes the order $G$ supertranslation. $v^{\mu}$ and $c^{\mu}$ are the 4-velocities and initial positions of the particle. $n^{\mu}=(1,n^{i})$ in the Cartesian coordinates where $n^{i}$ is the normal vector defined from the relation to the spatial parts of the Cartesian coordinates $x^{i}=rn^{i}$. The sign $\sum$ denotes the sum of contributions from all particles. For other notations, we would refer to [42]. Inserting the order $G$ metric into the effective stress tensor, we obtain, up to relevant order in $1/r$ expansion,
$T^{E}_{rr}=-\frac{C^{AB}C_{AB}}{2r^{4}}+{\cal O}(r^{-5}),\quad T^{E}_{ur}={\cal O}(r^{-4}),$
$T^{E}_{AB}=\frac{\gamma_{AB}C^{EF}N_{EF}}{2r}+{\cal O}(r^{-2}),\quad N_{AB}=\partial_{u}C_{AB},$
$T^{E}_{rA}=\frac{D_{A}(C^{AB}C_{AB})}{4r^{3}}+{\cal O}(r^{-4}),$ (11)
$T^{E}_{uA}=-\frac{D^{B}C_{AE}{N_{B}}^{E}+3C_{AB}D_{E}N^{BE}}{2r^{2}}+{\cal O}(r^{-3}),$
$T^{E}_{uu}=-\frac{N^{AB}N_{AB}+2C_{AB}\partial_{u}N^{AB}}{2r^{2}}+{\cal O}(r^{-3}).$
It is clear that $N_{AB}=0$ from the expression in (8), which simply confirms that there is no news function at order $G$. The components $T^{E}_{uu}$ and $T^{E}_{uA}$ from the effective stress tensor decay faster than the case of a pointlike particle at large distance, see, e.g., the asymptotic behaviors of the stress tensor in the Appendix of [46]. As will be detailed in the next pages, those components of the effective stress tensor control the evolution of the asymptotic charges from the self-part, the absence of which means that there is no source driving the next-to-leading order charges of the self-part. In the standard asymptotic analysis [5, 6], it is normally dealing with vacuum solution space, which only means that there is no local source closed to the null infinity. There can be matter fields inside the bulk of the spacetime. The non-trivial asymptotic charges simply indicate the information of the matter fields in the bulk. Since the effective stress tensor is defined locally, the linear and angular momentum from the self-part at the next-to-leading order should vanish with respect to the asymptotic behaviors of the effective stress tensor, which provides the main justification of our prescription.
We will prove that the order $G$ supertranslation freedom can be completely fixed by our prescription. We apply the asymptotic analysis to derive the generic order $G^{2}$ solutions with the effective stress tensor in series expansion near null infinity, where the initial data of the order $G^{2}$ metric will be fixed by the criteria of vanishing of the linear and angular momentum. We then compute the transformation laws of the linear and angular momentum under the order $G$ supertranslation. We show that the order $G$ supertranslation does not preserve the next-to-leading order angular momentum, hence, should be fixed by the vanishing condition of the angular momentum.
A curious point is that the supertranslation is only defined at null infinity but the PM system can be only solved in the bulk without touching any infinity data. The supertranslation dependence of the angular momentum loss from the bulk point of view can be better understood from the recently established amplitudes-based methods. The angular momentum loss at order $G^{2}$ is only relevant to the zero-frequency gravitons, see, e.g., [18, 19], which have clear connections to the null infinity supertranslation from the triangle equivalent relations [47, 48], see also [7].
3 Solution space at order $G^{2}$ and asymptotic charges
@@
We will solve the order $G^{2}$ self-part theory in series expansion near future null infinity. All of the metric components are assumed to be given in $1/r$ series expansion. For notational brevity, we suppress the label $S$ in the metric. We will adapt the Bondi gauge to the order $G^{2}$ theory where it is easy to read the following conditions $h_{rr}=h_{rA}=0$. The determinant condition in the Bondi gauge will induce constraints between the trace of $h_{AB}$ and the data at ${\cal O}(G)$. The precise condition is
@@
where $c_{AB}$ and $d_{AB}$ are traceless. This condition seems artificial. But the determinant condition indeed mixes data between different orders of the expansion in $G$. One can alternatively use the Newman-Unti (NU) gauge [49] (see also the application in linearized gravity [50]) where one can set the gauge condition $h_{ur}=0$. We have checked that transforming the solution in the NU gauge to the Bondi gauge will fix the subleading order of $h_{AB}$ as (12). The boundary conditions adapted from the Bondi gauge are $h_{ur}={\cal O}(r^{-2})$, $h_{uA}={\cal O}(1)$, and $h_{uu}={\cal O}(r^{-1})$. The effective stress tensor is conserved outside the trajectories of the pointlike bodies. Thus, we can use the well-known organization of the Einstein equation in the Bondi gauge [5, 6, 51, 50] to solve the order $G^{2}$ theory near null infinity. The four hypersurface equations $E_{r\alpha}=T^{E}_{r\alpha}$ will determine $h_{ur}$, $h_{uA}$, and $h_{uu}$. In $1/r$ series expansion, we obtain
@@
$h_{uA}=D^{B}c_{AB}+\frac{1}{r}\bigg{[}\frac{2}{3}(n_{A}+uD_{A}m_{b})-\frac{1}{16}D_{A}(C_{AB}C^{AB})\bigg{]}+{\cal O}(r^{-2}),$ (14)
@@
where $m_{b}$ and $n_{A}$ are integration constants from the differential equations of $r$. Two standard equations, the traceless part of $E_{AB}=T^{E}_{AB}$, will fix the evolution of the data at each order of $h_{AB}$ in the $1/r$ expansion except $c_{AB}$. In particular, we have $\partial_{u}d_{AB}=0$ from the leading order. The trace part of $E_{AB}=T^{E}_{AB}$ is trivial when the previous six equations of motion are satisfied. Three supplementary equations will fix the time evolution of the integration constants as
@@
$\partial_{u}m_{b}=\frac{1}{4}D_{A}D_{B}n^{AB}-\frac{1}{8}N_{AB}N^{AB},$ (17)
where $n_{AB}=\partial_{u}c_{AB}$. The results are consistent with the solution space of full Einstein theory near null infinity [51, 52]. Inserting the precise formula of $C_{AB}$ from the previous section, the evolution relations are reduced to
@@
The Bondi linear and angular momentum at the leading order in the expansion of $G$ coincide with the Arnowitt–Deser–Misner (ADM) definitions [42],
@@
The next-to-leading order contributions of the Bondi linear and angular momentum at null infinity are [53, 52]
@@
where $\text{d}\Omega$ is the integral measure on the celestial sphere and $Y^{(i)A}\frac{\partial}{\partial x^{A}}$ are three Killing vectors of the celestial sphere, which are related to the normal vector $n^{i}$ by $Y^{(i)A}=-\epsilon^{AB}D_{B}n^{i}$.
The vanishing of the next-to-leading order Bondi linear and angular momentum will fix $m_{b}$ and $n_{A}$ as
@@
$n_{A}=\frac{1}{4}\int\,\text{d}u\,\bigg{(}D^{B}D_{A}D^{E}c_{EB}-D^{2}D^{B}c_{BA}-uD_{A}D_{B}D_{E}n^{BE}\bigg{)}+......,$ (23)
where the ellipses denote $u$-independent total derivative terms which can not be fixed by the charge analysis and we have applied the identity [54, 31]
$-\frac{1}{4}Y^{(i)A}C_{AB}D_{E}C^{BE}=\frac{1}{4}D^{A}\bigg{[}Y^{(i)B}C_{AB}(D^{2}+2)\left[f(x^{A})-\bar{\beta}(x^{A})\right]\\ +\frac{1}{2}Y^{(i)}_{A}\big{(}(D^{2}+2)\left[f(x^{A})-\bar{\beta}(x^{A})\right]\big{)}^{2}\bigg{]},$ (24)
according to the expression of $C_{AB}$ in (8). Though we have fixed $m_{b}$ and $n_{A}$ from the asymptotic charges, the relevance to the order $G$ supertranslation is not yet uncovered. We will compute the transformation law of $m_{b}$ and $n_{A}$ under the order $G$ supertranslation and prove that a generic order $G$ supertranslation can not preserve the form of $n_{A}$ in (23).
4 Transformation laws under the order $G$ supertranslation
An order $G$ supertranslation connecting two solutions in the PM expansion up to order $G^{2}$ in the Bondi gauge can be effectively obtained by the coordinates transformation from $(u^{\prime},r^{\prime},{x^{\prime}}^{A})$ to $(u,r,x^{A})$. In $1/r$ expansion, we obtain up to relevant orders as
@@ 4 Transformation laws under the order $G$ supertranslation
$\quad\quad+2D_{A}TD^{A}T-\frac{1}{2}D^{2}TD^{2}T\bigg{)}+\frac{G^{2}}{6r^{2}}D^{AB}D_{A}D_{B}T+{\cal O}(G^{3})$ (26)
@@
$\quad-\gamma^{AB}\partial_{B}D^{C}TD_{C}T+\partial_{C}\gamma^{AB}D_{B}TD^{C}T\bigg{)}-\frac{G^{2}}{3r^{3}}D^{AB}D_{B}T+{\cal O}(G^{3}),$ (27)
where $T(x^{A})$ represents a supertranslation at order $G$, $\gamma=\sqrt{\det{\gamma_{AB}}}$, and $D_{AB}$ is the subleading order term of $H_{AB}$ in the $1/r$ expansion. We will not give the precise form of $D_{AB}$ as it will not affect the transformation law. Since our goal is to fix the order $G$ supertranslation, we did not include the order $G^{2}$ supertranslation. For computational simplicity, we choose the complex stereographic coordinates $A=(z,\bar{z})$ on the celestial sphere. The transformation laws of the Bondi mass aspect, angular momentum aspect, and asymptotic shear at order $G$ are obtained as
$\begin{split}&M_{b}={M^{\prime}}_{b},\quad\quad N_{A}={N^{\prime}}_{A},\\ &C_{AB}={C^{\prime}}_{AB}+(2D_{A}D_{B}-\gamma_{AB}D^{2})T,\end{split}$ (28)
where the primed quantities are functions in the old coordinates system. The results are in accordance with the fact that the leading order Bondi linear and angular momentum are irrelevant to the order $G$ supertranslation [42]. The transformation laws at order $G^{2}$ are
@@
$n_{A}={n^{\prime}}_{A}-D_{A}(T{M^{\prime}}_{b})-2{M^{\prime}}_{b}D_{A}T,$ (30)
@@
where we have used the expression of $C_{AB}$ in (8) to verify that $D^{\bar{z}^{\prime}}D^{\bar{z}^{\prime}}{C^{\prime}}_{\bar{z}^{\prime}\bar{z}^{\prime}}-D^{z^{\prime}}D^{z^{\prime}}{C^{\prime}}_{z^{\prime}z^{\prime}}=0$ when deriving (30). The obtained transformation laws are consistent with the finite transformation results in [55, 56, 57]. Clearly, the condition of preserving $n_{A}$ in the form of (23) in the new coordinates system will reduce the supertranslation to one translation along time direction $u$, considering that $M_{b}=-\sum{\frac{m}{(v_{\mu}n^{\mu})^{3}}}$ can not be a constant for arbitrary numbers of pointlike bodies. Even for the case of a single particle, $M_{b}$ is not necessary to be a constant as the boost charges of the supertranslated Schwarzschild black hole are generically non-vanishing [58]. The velocity of the particle essentially is a boost of the Schwarzschild solution [59].
5 Discussions
@@
In this paper, we apply the asymptotic charge analysis to prove that the order $G$ supertranslation should be ruled out in the PM expansion which can serve as the foundation of uniquely determining the angular momentum loss in gravitational scattering. It is worthwhile to point out that ruling out the order $G$ supertranslation only means that the supertranslation parameter is fixed, but not necessary to be zero. The precise formula of the order $G$ supertranslation parameter $\bar{\beta}(x^{A})$ can be obtained for a precise gravitational scattering, such as the gravitational Bremsstrahlung [60, 43, 12]. It is amusing to notice that once the order $G$ supertranslation is fixed, the complete order $G$ metric is fixed in the Bondi’s framework. The key point of our prescription is that we use part of the higher order solutions in the expansion of $G$ to fix the lower order supertranslation. It is interesting to check if such an idea can be valid at even higher orders in $G$ to fix the complete supertranslation freedom in the PM expansion.
As an ending remark, we will comment on the physical meaning of fixing the order $G$ supertranslation. The supertranslation dependence in the angular momentum is nothing but the fact that the angular momentum is observer dependent. This is a well known fact even in the classical mechanics. Hence, the results of the angular momentum loss in different literature can be different depending on the reference frames. Nevertheless, they must be connected by the order $G$ supertranslation. The key issue is that if one can associate the angular momentum and its loss to the proper observer and also specify their connections to the ADM value, e.g., the intrinsic gauge and canonical gauges discussed in [10] based on the far past conditions of the future null infinity [36], see also [61, 62] for earlier contributions along this direction. We conjecture that the precise form of the order $G$ supertranslation parameter $\bar{\beta}(x^{A})$ selected by our prescription should coincide with the canonical gauge at order $G$. Here, we offer a completely different prescription that is self-contained from the null infinity analysis.